### v2

# Evolutionary Reinforcement Learning via Cooperative Coevolution

Abstract
Recently, evolutionary reinforcement learning has obtained much attention in various domains. Maintaining a population of actors, evolutionary reinforcement learning utilises the collected experiences to improve the behaviour policy through efficient exploration. However, the poor scalability of genetic operators limits the efficiency of optimising high-dimensional neural networks. To address this issue, this paper proposes a novel cooperative coevolutionary reinforcement learning (CoERL) algorithm. Inspired by cooperative coevolution, CoERL periodically and adaptively decomposes the policy optimisation problem into multiple subproblems and evolves a population of neural networks for each of the subproblems. Instead of using genetic operators, CoERL directly searches for partial gradients to update the policy. Updating policy with partial gradients maintains consistency between the behaviour spaces of parents and offspring across generations. The experiences collected by the population are then used to improve the entire policy, which enhances the sampling efficiency. Experiments on six benchmark locomotion tasks demonstrate that CoERL outperforms seven state-of-the-art algorithms and baselines. Ablation study verifies the unique contribution of CoERL’s core ingredients.
1 Introduction
Reinforcement learning (RL) has made adequate advancements in various domains such as video games [20, 33], Go [29] and robotic control [10] with a competitive level beyond human. Evolutionary reinforcement learning (ERL) [11, 1] combines evolutionary computation and RL to solve sequential decision-making problems, capitalising on the benefits of evolution in exploration. ERL maintains a population of parameterised actors, called individuals, which interact with the paralleled environments to collect experiences. Individuals with higher reward-based fitness are more likely to be selected as the parents, to which mutation and crossover operators are applied to reproduce the next generation. Another RL agent called learner, learns from diverse experiences sampled by the maintained population of actors and substitutes the individual with the worst fitness in the population periodically, i.e., injecting its gradient information for better convergence.
However, ERL introduces a scalability problem, attributed to the genetic operators used for reproducing new individuals [43, 26, 14]. Minor perturbations in the parameter space, caused by genetic operators, lead to significant divergences in behaviour space [13, 14]. Consequently, an offspring might not inherit its parents’ behaviours by merely exchanging fragments of parameters. Thus, this inconsistency between parameter space and behaviour space does not align with the principles of evolution [42].
Lehman et al. [14] proposed a safe mutation operator based on the sensitivity of neural networks’ outputs, which slightly adjusts the parameters of neural networks to ensure consistency. State-based crossover [7], as well as distillation crossover and proximal mutation operator [2], improve the traditional genetic operators using backpropagation. However, those operators are gradient-based, which introduce in-negligible extra computational costs. Nevertheless, the works of [22] and [39] directly searched for the parameters of neural networks using the cooperative coevolution, in which the policy optimisation problem is decomposed into multiple low-dimensional subproblems. However, the experiences of optimising subproblems are not fully utilised, resulting in inefficient training [11].
To address the scalability issue, we propose a novel cooperative coevolutionary reinforcement learning (CoERL) algorithm. The optimisation of a high-dimensional neural network is decomposed into multiple low-dimensional subproblems periodically and adaptively by cooperative coevolution. A population is resampled for each subproblem, in which each individual maintains the same neural network architecture. CoERL directly searches for partial gradients and updates the entire policy. The partial gradient searched via the population guides a proximal update on the subproblem, diminishing inconsistency between the behaviour spaces of parents and offspring. Experiences collected during evolution are then used to enhance RL for better sampling efficiency.
Our main contributions are summarised as follows:
  • This paper proposes CoERL, a novel cooperative coevolutionary reinforcement learning algorithm to address the scalability issue of ERL, which also improves sample efficiency during training.
  • This paper proposes a partially updating strategy for policy improvement based on cooperative coevolution, which costs less and maintains the consistency between the behaviour spaces of parents and offspring.
  • Experimental results on six benchmark locomotion tasks demonstrate the comparable performance of CoERL, compared with seven state-of-the-art algorithms and baselines. Ablation study shows the unique contributions of CoERL’s core ingredients.
2 Background
In this section, we introduce some preliminary, cooperative coevolution and recent progress of ERL.
2.1 Preliminary: Markov decision process
Markov decision process (MDP) [30] is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma)$, where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$ is the transition probability function, and $\gamma$ is the discount factor. A policy $\pi$ is a mapping from states to probability distributions for acting, where $\pi(a_{t}|s_{t})$ is the probability of taking action $a_{t}$ in state $s_{t}$ at time $t$. The goal of the MDP is to optimise a parameterised policy $\pi_{\theta}$ that maximises the discounted cumulative reward, formulated as:
$\max_{\theta}~{}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(s_{t},a_{t},s_{t+1})],$ (1)
where $s_{0}$ is an initial state. $\tau\sim\pi_{\theta}$ denotes a trajectory $(s_{0},a_{0},s_{1},a_{1},\dots,s_{t},a_{t},s_{t+1})$ sampled from $\pi_{\theta}$.
2.2 Cooperative coevolution
Cooperative coevolution (CC) simulates the interactions of species living in an ecosphere [18].  Potter and De Jong [24] proposed the first CC algorithm, called cooperative coevolutionary genetic algorithm (CCGA). CCGA decomposes an optimisation problem into multiple subproblems. A population is initialised for each subproblem with a context vector for fitness evaluations. For example, given an optimisation problem with $n$ variables, the fitness of a partial solution $x_{ij}$, i.e., the $j$-th individual in the population of $i$-th subproblem, can be evaluated by replacing the variable $b_{j}$ in the context vector $\bm{b}$. Each variable in $\bm{b}$ is updated with the best individual for the corresponding subproblem.
Problem decomposition referring to the grouping of variables is critical in CC [18]. Incorrect grouping leads to a local optimum, which is limited by the decomposed problem itself [23, 38]. Intuitively, static grouping and random grouping [24] are easy to implement and require no additional computational budget or knowledge. However, the effectiveness of the general strategies is unclear and requires some presets. Interaction-based grouping [16] and landscape-aware grouping [37] decompose the problem according to the characteristics of variables such as correlation and landscape. However, additional costs for fitness evaluations are required.
Another issue concerns the evaluation of individuals within a subproblem’s population, also known as the credit assignment problem, since individuals in a population represent only part of the entire solution. Collaboration for fitness evaluation with other subproblems’ populations is required. A global context vector can be constructed by collaborators (i.e., individuals) selected from populations of subproblems. Popular selection methods, such as single best collaborator [24], elite collaborator [8], and random collaborator selection, are typically chosen based on the specific domain [18].
Cooperative coevolution, served as a useful technique for solving high-dimensional black-box optimisation problems, is expected to benefit the policy optimisation. In this paper, we apply random grouping and single best collaboration for low cost and simplicity.
Figure 1: Inconsistency phenomenon between the behaviour spaces of parents and offspring by applying one-point crossover to exchange partially the parameters of Actor 1 (red) and Actor 2 (green). Feature maps show the behaviour spaces decompressed by t-SNE [32].
2.3 Evolutionary reinforcement learning
Evolutionary algorithms (EAs) have been successfully applied to solve RL problems, where neuroevolution and evolution-guided policy gradient are two main trends in this area [21, 43, 36, 28, 1].
2.3.1 Neuroevolution
Neuroevolution directly optimises the parameters or architectures of neural networks without considering the underlying mechanism like gradients, making it suitable for evolving reinforcement learning policy and tackling real-world problems [17, 4, 15, 6, 40]. Fitness metrics formulated on rewards are used to select good individuals for reproducing the next generation by mutation and crossover. Hence, neuroevolution can intuitively handle complicated rewards such as non-convex and non-differentiable cases [26, 4]. Population-based training also encourages exploration. However, Nesterov and Spokoiny [22] indicated that EAs scale poorly with the increase of parameters.
Gong et al. [9] combined coevolution and backpropagation to solve classification tasks, while the framework is limited by traditional reproduction and specific domains. Yang et al. [39] proposed a CC framework based on negatively correlated search (NCS) to tackle this issue, which directly searches for the parameters of a neural network. However, it only utilises the final accumulative reward as fitness, resulting in a waste of experiences collected during evolution.
Besides, traditional genetic operators limit EA’s capability of addressing large-scale optimisation problems. The direct operations on the parameter space lead to an inconsistency between the behaviour spaces of parents and offspring, where parental behaviours may be typically forgotten [14, 7, 2]. Figure 1 shows an example of applying the one-point crossover, in which behaviour spaces of actors are decompressed by t-SNE [32]. It is easy to see from Figure 1 that directly exchanging the parameter fragments of the parents leads to a forgetting phenomenon in the behaviour space of the offspring.
2.3.2 Evolution-guided policy gradient
Figure 2: Diagram of CoERL. The CoERL algorithm begins by decomposing the policy optimisation problem parameterised by a neural network into multiple subproblems and searching for partial gradients to update the policy. Subsequently, the explored experiences are gathered to further refine the policy outcome using the MDP-based RL.
Evolution-guided policy gradient combines neuroevolution and MDP-based RL [11]. In the EA loop, actors, also called individuals, interact with the environments and are evolved by genetic operators. Experiences collected during the evaluations, i.e., transitions, are stored in a replay buffer and leveraged in the RL loop. Khadka et al. [12] proposed a portfolio of policies and dynamically allocated computational resources to train different policies, extending ERL. CEM-RL [25] leveraged cross-entropy method (CEM) to delayed deep deterministic policy gradient (TD3). Without using genetic operators, CEM-RL samples actors from the policy distribution estimated in previous generations. Gangwani and Peng [7] applied imitation learning to the crossover operator, which use trajectories of two parents to train an offspring. Bodnar et al. [2] improved the mutation and crossover by using local replay memory and the critic of the RL learner, namely proximal distilled evolutionary reinforcement learning (PDERL). New offsprings are produced by exchanging the local replay buffer or perturbation according to the sensitivity of the actions in PDERL [2]. Genetic algorithm has been combined with RL in the work of [19], where individuals are generated by the noisy mutation only. To tackle the computationally expensive evaluation, Wang et al. [34] extended the surrogate model to ERL, which estimates the fitness of each individual based on the approximated value function. Zhu et al. [44] addressed the exploration and exploitation issue by maintaining actor and critic populations.
Although ERL inherits the advantages of neuroevolution, enabling it to address the temporal credit assignment problem with sparse rewards, it faces the challenge of poor scalability with EAs. Recent efforts have primarily concentrated on enhancing genetic operators, but at the expense of increased computational costs.
3 Cooperative coevolutionary reinforcement learning
Aiming at addressing the poor scalability of ERL, we propose a novel cooperative coevolutionary reinforcement learning (CoERL) algorithm. Pseudo-code and diagram of CoERL are provided in Algorithm 1 and Figure 2, respectively.
Algorithm 1 CoERL.
1:The number of generations $T$, population size $\mu$, noise strength $\sigma$, number of subproblems $m$, learning rate $\alpha$, temperature coefficient $\alpha_{s}$
2:Policy $\pi_{\theta}$
3:Initialise policy $\pi_{\theta}$, two critics $\hat{Q}_{\phi_{1}}$ and $\hat{Q}_{\phi_{2}}$
4:Initialise reward buffer $\mathcal{B}_{\mathcal{R}}$
5:for $n=1$ to $T$ do
6:     Decompose the policy optimisation problem into $m$
     sub-problems by grouping $|\theta|$-d parameters to $m$ sets
     of parameter indices $\mathcal{I}_{1},\cdots,\mathcal{I}_{m}$
7:     for $j=1$ to $m$ do
8:         Sample $\epsilon_{1},\cdots,\epsilon_{\mu}\sim\mathcal{N}^{|\mathcal{I}_{j}|}(0,I)$
9:         Reproduce a population $\{\psi_{1},\cdots,\psi_{\mu}\}$ for sub-
         problem $\mathcal{I}_{j}$ by $\psi_{i}\leftarrow\theta^{\langle\mathcal{I}_{j}\rangle}+\epsilon_{i}$, for $i=1,\cdots,\mu$
         $\triangleright$ $\theta^{\langle\mathcal{I}_{j}\rangle}$ denotes $\theta$’s elements at indices $\mathcal{I}_{j}$
10:         for $i=1$ to $\mu$ do
11:              $J^{\pi_{\psi_{i}}}_{\mathcal{R}},\tau^{\pi_{\psi_{i}}}=\text{Evaluate}(\pi_{\psi_{i}})$
12:              Store $\tau^{\pi_{\psi_{i}}}$ in $\mathcal{B}_{\mathcal{R}}$
13:              Assign fitness $f_{i}=J^{\pi_{\psi_{i}}}_{\mathcal{R}}$
14:         end for
15:         $\theta^{\langle\mathcal{I}_{j}\rangle}\leftarrow\theta^{\langle\mathcal{I}_{j}\rangle}+\alpha\frac{1}{\mu\sigma}\sum^{\mu}_{i=1}f_{i}\epsilon_{i}$triangleright Eq. (5)
         $\triangleright$ Update parameters of policy $\pi_{\theta}$ indexed by $\mathcal{I}_{j}$
16:     end for
17:     $\triangleright$ Policy gradient with $\mathcal{B}_{\mathcal{R}}$:
18:     Randomly sample a minibatch $\mathcal{B}$ of transitions
     $\mathcal{T}=\langle s,a,s^{\prime},r\rangle$ from $\mathcal{B}_{R}$
19:     Compute target
     $y=r-\gamma(\min\limits_{j=1,2}\hat{Q}_{\phi_{j}}(s^{\prime},\tilde{a}^{\prime})-\alpha_{s}\log{\pi_{\theta}(\tilde{a}^{\prime}|s^{\prime})})$,
     where $\tilde{a}^{\prime}\sim\pi_{\theta}(\cdot|s^{\prime})$
20:     Update critics with
     $\nabla_{\phi_{j}}\frac{1}{|B_{R}|}\sum\limits_{\mathcal{T}\in\mathcal{B}}(y-\hat{Q}_{\phi_{j}}(s,a))^{2}$ for $j=1,2$
21:     Update actor with
     $\nabla_{\theta}\frac{1}{|B_{R}|}\sum\limits_{\mathcal{T}\in\mathcal{B}}(\min\limits_{j=1,2}\hat{Q}_{\phi_{j}}(s,\tilde{a}_{\theta})-\alpha_{s}\log\pi_{\theta}(\tilde{a}_{\theta}|s))$,
     where $\tilde{a}_{\theta}$ is sampled from $\pi_{\theta}(\cdot|s)$ via the reparame-
     terisation trick
22:end for
3.1 Collaboration between cooperative coevolution and reinforcement learning
CoERL maintains collaborative loops between CC and RL. In the CC loop, a neural network (a policy) is converted into a position-fixed vector with real numbers. The policy optimisation problem, is divided into multiple subproblems, in which parameters of the neural network are grouped. The decomposed subproblems have no intersection, while their union collectively constitute the entire parameter space. Given a policy parameterised by $\theta$, it ensures that $\theta^{\langle\mathcal{I}_{i}\rangle}\cap\theta^{\langle\mathcal{I}_{j}\rangle}=\varnothing,\forall i\neq j,1\leq i\leq m,1\leq j\leq m$ and $\theta^{\langle\mathcal{I}_{1}\rangle}\cup\theta^{\langle\mathcal{I}_{2}\rangle}\cup\cdots\cup\theta^{\langle\mathcal{I}_{m}\rangle}=\theta$, where $m$ is the number of subproblems and $\langle\mathcal{I}_{i}\rangle$ denotes the parameter indices of subproblem $j$. To optimise a subproblem, a population is maintained and generated using perturbation methods instead of traditional genetic operators. Practically, each individual in the population shares the same neural network structure with the freezing quotient set of the subproblem. Only the variables in the subproblem, which form part of the neural network, are perturbed by the noises from the distribution.
Individuals consistently interact with the environment. The cumulative reward obtained by the individual is treated as its fitness. Then, a number of noised individuals and their corresponding fitness are used to update the entire policy. Notably, subproblems are not optimised separately. The policy optimised according to the preceding subproblem is regarded as the complementary base of the subsequent subproblem. Thus, connections between different subproblems are built.
After updating the policy within each subproblem, CoERL proceeds to optimise the policy via MDP-based RL. Experiences such as transitions produced by the individuals during the cooperative coevolution loop are collected in a replay buffer. Batches sampled from the replay buffer are then used to update the policy via policy gradient, which fully utilises the temporal information. Sections 3.2 and 3.3 detail the two loops, respectively.
3.2 Partially updating via cooperative coevolution
Given a policy $\pi_{\theta}$ parameterised by $\theta$, CoERL decomposes the policy optimisation problem into $m$ subproblems in the CC loop. The subproblem $j$ of policy optimisation, $\theta^{\langle\mathcal{I}_{j}\rangle}$ indexed by parameters indices $\langle\mathcal{I}_{j}\rangle$, and the quotient $\hat{\theta}^{\langle\mathcal{I}_{j}\rangle}=\theta/\theta^{\langle\mathcal{I}_{j}\rangle}$ constitute the entire parameter space, i.e., $\theta=\left[\theta^{\langle\mathcal{I}_{j}\rangle}:\hat{\theta}^{\langle\mathcal{I}_{j}\rangle}\right]$. Then, for each subproblem $j$, a corresponding actor $\pi_{\theta^{\langle\mathcal{I}_{j}\rangle}}$ is constructed by freezing the quotient part. CoERL partially updates the policy by iteratively optimising each subproblem. A population is first sampled via a specific distribution $P^{\langle\mathcal{I}_{j}\rangle}$. Each individual is an actor $\pi_{\theta^{\langle\mathcal{I}_{j}\rangle}}$, where $\theta^{\langle\mathcal{I}_{j}\rangle}\sim P^{\langle\mathcal{I}_{j}\rangle}$. All individuals in the population of the subproblem share the same quotient part $\hat{\theta}^{\langle\mathcal{I}_{j}\rangle}$. Then, the fitness of each individual $f(\pi_{\theta^{\langle\mathcal{I}_{j}\rangle}})$ is determined by the cumulative reward. The expected fitness under an arbitrary distribution is formulated as Eq. (2):
[…]
where $p(\theta^{\langle\mathcal{I}_{j}\rangle})$ denotes the density. Then, we can write the gradient form using the “log-likelihood trick". The estimation of the gradient by maintaining a population with size $\mu$ is shown as follows:
$\nabla_{\theta^{\langle\mathcal{I}_{j}\rangle}}J(\theta^{\langle\mathcal{I}_{j}\rangle})=\mathbb{E}_{\theta^{\langle\mathcal{I}_{j}\rangle}}\left[f(\pi_{\theta^{\langle\mathcal{I}_{j}\rangle}})\nabla_{\theta}\log(p(\theta^{\langle\mathcal{I}_{j}\rangle}))\right].$ (3)
In the case of factored Gaussian distribution with the deviation $\sigma$, we can set $\theta^{\langle\mathcal{I}_{j}\rangle}$ as the mean vector. Then, the expected fitness of Eq. (2) is rewritten as follows:
$\mathbb{E}_{\theta^{\langle\mathcal{I}_{j}\rangle}}[f(\pi_{\theta^{\langle\mathcal{I}_{j}\rangle}})]=\mathbb{E}_{\epsilon\sim\mathcal{N}^{\langle\mathcal{I}_{j}\rangle}(0,I)}[f(\pi_{\theta^{\langle\mathcal{I}_{j}\rangle}+\sigma\epsilon})].$ (4)